# A Benchmark for Data Management in Microservices

### v2

@@ Abstract.
Microservice architectures emerged as a popular architecture for designing scalable distributed applications. Although microservices have been extensively employed in industry settings for over a decade, there is little understanding of the data management challenges that arise in these applications. As a result, it is difficult to advance data system technologies for supporting microservice applications. To fill this gap, we present Online Marketplace, a microservice benchmark that incorporates core data management challenges that existing benchmarks have not sufficiently addressed. These challenges include transaction processing, query processing, event processing, constraint enforcement, and data replication. We have defined criteria for various data management issues to enable proper comparison across data systems and platforms.
After specifying the benchmark, we present the challenges we faced in creating workloads that accurately reflect the dynamic state of the microservices. We also discuss issues that we encountered when implementing Online Marketplace in state-of-the-art data platforms and meeting the criteria. Our evaluation demonstrates that the benchmark is a valuable tool for testing important properties sought by microservice practitioners. As a result, our proposed benchmark will facilitate the design of future data systems to meet the expectations of microservice practitioners.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/diku-dk/EventBenchmark.
@@ 1. Introduction
Despite the benefits of the decoupled design, a recent study (Laigner et al., 2021) demonstrates that practitioners encounter several challenges when trying to meet data management requirements in this architecture, including distributed transaction processing, data replication, consistent data and event querying and processing, and enforcing data integrity constraints. In short, microservices are designed to function independently, but in practice, they often rely on each other’s data and functionality to complete a workflow. As a result, it is essential to benchmark microservices in a way that accurately reflects the needs of practitioners. Unfortunately, existing microservice benchmarks do not fully capture these real-world requirements (Laigner et al., 2021; Gan et al., 2019; Zhou et al., 2021). For example, DeathStarBench (Gan et al., 2019) does not consider event-driven architecture, nor does it specify what data invariants and transactional guarantees are necessary.
@@
Online Marketplace can benchmark various systems, platforms, and frameworks for developing data-intensive microservices. We verify the applicability of our benchmark by implementing five versions of Online Marketplace in two different platforms, Orleans and Statefun, which are designed for event-driven, distributed stateful applications. We conducted experiments to measure how these competing platforms perform under different workload scenarios. During the implementation of Online Marketplace on these state-of-the-art platforms, we encountered several limitations that prevented us from fulfilling some of the data management functionalities and criteria defined by Online Marketplace. Meanwhile, our results show that our benchmark can effectively stress the performance of the platforms and reveal performance and functionality issues. As a result, our proposed benchmark will facilitate the design of futuristic data systems that can fully meet the expectations of microservice practitioners, who are an important part of the database user community.
2. Motivation
Data Management Challenges in Practice. Despite more than a decade-long employment in the industry, little work in data management research has incorporated the challenges brought about by data-intensive applications following the microservices architectural style. Based on the findings of a recent study that empirically investigated the data management challenges in microservices (Laigner et al., 2021) and public reports (Gluck, 2022; Yang et al., 2021; Singla and Wu, 2020; Chornyi, 2018; Ferreira, 2019; Ferreira and Wible, 2017; Sarma, 2020; Andreakis et al., 2019; Khisti, 2018; Chew, 2019; Silnitsky, 2022a, b; Stenberg, 2019; Mayerhofer, 2018; Oliveira, 2017; Conger, 2021; Viennot et al., 2015) of companies with large-scale deployments, we summarize the core data management challenges affecting practitioners as follows:
@@ 2. Motivation
Problems of Existing Benchmarks. Existing benchmarks like DeathStarBench (Gan et al., 2019) and TrainTicket (Zhou et al., 2018) do not fully capture these real-world challenges. They do not take into consideration event-driven architecture; thus, challenges related to correct event processing are missing. Furthermore, they do not consider any requirements on data invariants, transactional guarantees, data replication or data querying.
Therefore, these benchmarks cannot be used to benchmark the overhead of addressing the correlated data management challenges, which includes the overhead of all-or-nothing atomicity, concurrency control, consistent query processing, consistent data replication, and correct event processing. Consequently, they cannot be used to facilitate the development of new data systems to meet the real-world data management challenges of microservices.
@@ 3.1. Design Goals
Generalizability. Many other popular application domains can be mapped to an online marketplace. These include multi-tenant e-commerce platforms, airline retailing (IATA, [n.d.]) (e.g., hotels, travel agencies, and taxi and insurance companies as sellers), social media marketplaces, delivery service (e.g., connecting individuals to a myriad of businesses, like restaurants), digital investment (e.g., funds from different banks as products), and homestay platforms. It is noteworthy that many of these applications report data management challenges that we address in this work:
  • Event ordering and data invariants (Ferreira and Wible, 2017; Sarma, 2020; Yang et al., 2021; Singla and Wu, 2020; Chew, 2019; Silnitsky, 2022a, b)
  • Transactional consistency across microservices (Gluck, 2022; Khisti, 2018; Chornyi, 2018; Ferreira, 2019; Stenberg, 2019)
  • Consistent data querying  (Mayerhofer, 2018; Oliveira, 2017; Conger, 2021)
  • Data replication correctness (Viennot et al., 2015; Sarma, 2020; Andreakis et al., 2019)
@@ 3.3. Microservices
[Partial Data model] Refer to caption [Events] Refer to caption
@@
Traditionally, a microservice-based application is often made up of independent services, each deployed in a container on an OS-level virtualized platform such as Docker (Laigner et al., 2021). However, the emergence of programming frameworks for designing distributed applications, such as Orleans (Bykov et al., 2010) and Statefun (Statefun, 2023c), provides programming models that allow microservices to be implemented using higher-level abstractions, such as actors and functions, respectively.
To allow this benchmark to support the gamut of microservice implementations, we withdraw ourselves from defining an architectural blueprint, but rather, we focus on describing the expected independent components that must compose the benchmark application. In this sense, in the following paragraphs, we take advantage of Figure 1 to clarify the interactions, APIs, and the data model prescribed for each microservice. It is worth noting that we use the term ”events” as a communication abstraction for microservices, but these can also be framed as any message payload asynchronously delivered to a microservice. Besides, the event identifiers in Figure 1(b) do not imply a particular order in which the microservices exchange events.
Furthermore, while we specify the microservices’ state and the queries using the relational model, they can also be specified through other data and query models as long as the same functionalities can be achieved.
Cart. The Cart microservice allows the customer to manage the products that ought to compose a checkout order. It does so through the following APIs:
  • AddItem. Add a product with an associated quantity and possible discount to a customer’s cart.
  • RemoveItem. Remove a product from a customer’s cart.
  • Checkout. Submit a cart for checkout and seal the cart.
To manage a customer’s cart, two relations are used: carts, to track the status of a customer session, and $cart\_items$, to store the items that customers are willing to buy. Moreover, the $replica\_products$ relation represents a partial replica of the product table belonging to the Product microservice. The table is only updated through ProductUpdate events received from the Product microservice (represented by a dotted blue arrow in Figure 1(a)). The replication semantics are discussed in Section 4.3.
When a checkout request is received, Cart ensures the correctness of the customer’s cart by matching the cart items to the replicated product information and looking for products that had their price updated. In this case, the price is updated and the difference (if positive) between the new and the old price becomes a discount in the cart item. The customer’s cart items are then assembled into a ReserveInventory event and published for asynchronous processing. Once the event is published, the cart is sealed, allowing the customer to initiate a new cart and execute a series of cart operations.
@@
  • GetProduct. Retrieve a product based on a seller and a product identifier.
@@
Stock. The Stock microservice manages the inventory through a single relation: $stock\_item$. The following API is provided:
  • GetStockItem. Retrieve the information of a stock item.
  • AddStockItem. Add a stock item with an associated quantity available.
To update the inventory, the Stock microservice processes four events: ReserveInventory, PaymentConfirmed, PaymentFailed, and ProductDelete. Every inbound event leads to updates in one or more $stock\_item$’ tuples (depending on the number of items present in the checkout).
@@
Upon receiving a StockConfirmed event, the amount to charge the customer is calculated (including freight and discounts) and an invoice number is generated, resulting in a InvoiceIssued event. During this processing, all four relations have tuples created (in case it is the first customer order, otherwise $customer\_order$ tuple is updated).
Afterwards, there are two moments where a customer’s order status is updated: after a payment and after shipment updates. Both trigger writes to order and $order\_history$ relations. In this case, the order of event processing here plays a role on maintaining the notion of time progress for individual orders.
Payment. The Payment microservice is responsible for managing payment data. Two relations are used to track an orders’ payment data: $payment$ and $card\_payment$.
Payment processing starts through processing an InvoiceIssued event. Every credit applied to the order, namely, discounts and the payment option chosen by the customer at the time of the checkout (credit/debit card or a bank slip) are stored in $payment$ relations and in case of a card payment, a $card\_payment$ tuple is also inserted.
@@
The relation $order\_entry$ represents an order item, but from the perspective of a seller. In this sense, amounts related to the item (i.e., discount, freight, total) are calculated. Basically every order (derived from the InvoiceIssued payload) is transformed into $N$ $order\_entry$ tuples, where N is the number of items corresponding to a seller. $order\_entry\_details$ relation contains data agnostic to sellers, such as payment method used and total order amount. Besides, the relation $order\_entry$ is updated through processing the following events: PaymentConfirmed, PaymentFailed, ShipmentNotification, and DeliveryNotification.
@@ 3.4.1. Business Transactions
To realize the application scenario, we describe four business transactions that reflect different complexities in terms of number of involved microservices and number of events processed. In Section 4, we discuss how transactions possess different properties.
Customer Checkout. It starts in Cart microservice through processing a Checkout request. It involves Cart, Stock, Order, Payment, Shipment, Seller, and Customer microservices. A success path involves the following events: E2 - E3 - E5 - E6 - E8. There are two error cases: (i) when payment fails, then from E5 we have E7 as the last; and when no single item from the cart is available in stock: E2 - E4. Although not shown in the Figure 1 due to space constraints, a success path also involves E6 being processed by Order microservice.
Price Update. To enable the partial replication of products in the Cart, upon processing a UpdatePrice request and updating its private state accordingly, the Product microservice generates E0 and sends it to the Cart microservice. By processing E0, the Cart microservice applies the new price to its corresponding product replica.
Product Delete. To simulate making a product unavailable to customers, we pick a seller and a corresponding product (both from a distribution) and we set the product as disabled. To maintain the total number of products, thus avoiding anomalies in the distribution, we replace the deleted product with another one.
This operation is carried out through processing a DeleteProduct request in Product microservice. Upon updating its private state accordingly, Product generates E1 and sends it to the Stock microservice.
Update Delivery. To simulate the delivery of goods, we pick the first 10 sellers with uncompleted orders (i.e., at least one package has not been delivered yet) on chronological order and we set their respective oldest order’s packages as delivered.
The idea is that packages are progressively delivered as more update delivery transactions are submitted to the system. For maintenance of statistics about customer and orders, the following events are generated along with the transaction: E9 is generated by Shipment for every package delivered and sent to the Seller and Customer microservices; and E8 is generated by Shipment when all packages of a shipment are delivered and sent to the Order microservice.

### v4

# Online Marketplace: A Benchmark for Data Management in Microservices

@@ Abstract.
Microservice architectures have become a popular approach for designing scalable distributed applications. Despite their extensive use in industrial settings for over a decade, there is limited understanding of the data management challenges that arise in these applications. Consequently, it has been difficult to advance data system technologies that effectively support microservice applications. To fill this gap, we present Online Marketplace, a microservice benchmark that highlights core data management challenges that existing benchmarks fail to address. These challenges include transaction processing, query processing, event processing, constraint enforcement, and data replication. We have defined criteria for various data management issues to enable proper comparison across data systems and platforms. Through case studies with state-of-the-art data platforms, we discuss the issues encountered while implementing and meeting Online Marketplace’s criteria. By capturing the overhead of meeting the key data management requirements that are overlooked by existing benchmarks, we gain actionable insights into the experimental platforms. This highlights the significance of Online Marketplace in advancing future data systems to meet the needs of microservice practitioners.
microservices, data management, benchmark, online marketplace
copyright: rightsretainedjournal: PACMMODjournalyear: 2025journalvolume: 3journalnumber: 1 (SIGMOD)article: 3publicationmonth: 2doi: 10.1145/3709653ccs: Information systems Data management systemsccs: Computer systems organization Distributed architectures
@@ 1. Introduction
This industrial popularity has prompted cloud providers to offer rich features targeting microservice deployments (ms_, 2024a, d, c, b), such as specialized container-based technologies to deploy and scale microservices, message brokers to allow for loose-coupled microservices, multi-tenant DBMS technologies to meet the isolation properties of microservices, and application frameworks and side-car technologies to speed up deployment, code evolution and maintenance of microservices.
Despite the benefits of the decoupled design, a recent study (Laigner et al., 2021) demonstrates that practitioners encounter several challenges when trying to meet data management requirements in this architecture, including distributed transaction processing, data replication, consistent data and event querying and processing, and enforcing data integrity constraints. In short, microservices are designed to function independently, but in practice, they often rely on each other’s data and functionality to complete a workflow. As a result, it is essential to benchmark microservices in a way that accurately reflects the needs of practitioners. Unfortunately, existing microservice benchmarks do not fully capture these real-world requirements (Laigner et al., 2021; Gan et al., 2019; Zhou et al., 2021). For example, DeathStarBench (Gan et al., 2019) and TrainTicket (Zhou et al., 2018) do not consider event processing, nor do they specify what data invariants and transactional guarantees are necessary. The absence of a comprehensive benchmark for measuring the performance of data management tasks in microservices leads to the development of custom benchmarks that fail to reflect the complexities of data management in real-world microservices (Burckhardt et al., 2022).
@@
We verify the applicability of our benchmark by implementing five versions of Online Marketplace in three state-of-the-art data platforms, Orleans, Statefun, and a composite solution. Orleans and Statefun are designed for event-driven, distributed stateful applications. The composite solution reflects a usual architecture seek in practice (Laigner et al., 2021; Kleppmann et al., 2019). While implementing Online Marketplace on these platforms, we encountered several limitations that prevented us from fulfilling some of the data management functionalities and criteria sought in practice. We conducted experiments to measure how these competing platforms perform under different workloads and observed their design and architectural implications in performance. Our results show that our benchmark can effectively stress the performance of the platforms and reveal performance and functionality issues. Meanwhile, through Online Marketplace, we derive actionable insights to foment the design of futuristic data systems that will meet the expectations of microservice practitioners, who are an important part of the database user community.
By providing a concrete example for database researchers to understand the unmet needs of data management in microservices, the benchmark serves as a testbed for experimenting with novel algorithms, approaches, techniques, and programming models. For instance, a recent study (Liu et al., 2024) aims to fulfill the data management criteria outlined here to enhance state management features in actor systems for cloud-native applications. Furthermore, Online Marketplace is a benchmark that not only benefits the real-world microservice applications that served as inspiration for our design (Gluck, 2022; Yang et al., 2021; Singla and Wu, 2020; Chornyi, 2018; Ferreira, 2019; Ferreira and Wible, 2017; Sarma, 2020; Andreakis et al., 2019; Khisti, 2018; Chew, 2019; Silnitsky, 2022a, b; Stenberg, 2019; Mayerhofer, 2018; Oliveira, 2017; Conger, 2021; Viennot et al., 2015) but can also extend beyond existing microservices applications. For instance, besides microservices, there are trends of other architectural styles and programming models offering trade-offs between complexities in data management and the degrees of coupling. These include modular monolithic architectures, FaaS, Actors, etc. As demonstrated in our case studies (§$\S$ 7), Online Marketplace is well positioned to serve as a testbed for these trending cloud application architectures and programming models. The artifacts used in this work are available online (Laigner, 2023) for reproducibility and extension.
2. Background
Data Management Challenges. Despite more than a decade-long employment in the industry, little work in data management research has incorporated the challenges brought about by data-intensive applications following the microservices architectural style. Based on the findings of a recent study that empirically investigated the data management challenges in microservices (Laigner et al., 2021) and public reports (Gluck, 2022; Yang et al., 2021; Singla and Wu, 2020; Chornyi, 2018; Ferreira, 2019; Ferreira and Wible, 2017; Sarma, 2020; Andreakis et al., 2019; Khisti, 2018; Chew, 2019; Silnitsky, 2022a, b; Stenberg, 2019; Mayerhofer, 2018; Oliveira, 2017; Conger, 2021; Viennot et al., 2015) of companies with large-scale deployments, we summarize the core data management challenges as follows:
@@ 2. Background
Table 1. Comparison of microservice benchmarks
Requirement Criteria DeathStar & TrainTicket Online Marketplace
Functional Decomposition Isolation of Resources Yes Yes
Asynchronous Events Event Processing Order No Yes
Event Processing and Delivery Guarantees No Yes
Distributed Transactions All-or-nothing Atomicity; Isolation Levels Allowed No Yes
Data Invariants Data Invariant Enforcement No Yes
Data Replication Data Caching and Replication Consistency No Yes
Query Processing Query Processing Consistency No Yes
Problems of Existing Benchmarks. Existing microservice benchmarks like DeathStarBench (Gan et al., 2019) and TrainTicket (Zhou et al., 2018) do not fully capture these real-world challenges, as shown in Table 1. DeathStar  (Gan et al., 2019) was created to investigate the effects of microservice architectures on hardware and software in system stacks, particularly network and operating systems. TrainTicket focuses on replicating industrial faults in microservices, supporting researchers on investigating fault analysis and debugging. However, they do not take into consideration asynchronous events; thus, challenges related to correct event processing are missing. Furthermore, they do not consider any requirements on data invariants, transactional guarantees, data replication or data querying. Therefore, they target problems that are either oblivious to data management or position data management as not a primary concern.
Similar to microservices, Object-Oriented DBMSs (OODBMS) also adopt the modularization principle, such as encapsulating state and operations into objects within the database. Although, in principle, a microservice can be mapped to an object in OODBMS, we find that OODBMS benchmarks cannot meet the needs that we target. Benchmarks like Cattell, HyperModel, OO7, Justitia, and OCB aim for modeling object relationships in the context of engineering applications and evaluating object clustering algorithms (Darmont and Schneider, 2002). All these fail to capture the data management requirements in Table 1.
TPC-W (Council, 2005) is a transactional benchmark that models the core aspects of user experience on an e-commerce website, such as browsing pages, checking out books, and searching for keywords. TPC-C (Council, 2010) was designed to reflect the transaction processing of a wholesale supplier. YCSB (Cooper et al., 2010) models transactional workloads in the cloud that are not necessarily executed under ACID semantics. All of them assume a traditional monolithic application architecture. Our benchmark models a modern microservice-oriented application and the complex interplay of their components through events, differing substantially from these benchmarks regarding architectural style and data management requirements.
Therefore, existing benchmarks cannot be used to benchmark the overhead of addressing the correlated data management challenges, which includes the overhead of all-or-nothing atomicity, concurrency control, consistent query processing, consistent data replication, and correct event processing. Consequently, they cannot be used to facilitate the development of new data systems to meet the real-world data management challenges of microservices.
@@ 3.1. Design Goals
Generalizability. Many other popular application domains can be mapped to an online marketplace. These include multi-tenant e-commerce platforms, airline retailing (IATA, [n.d.]) (e.g., hotels, travel agencies, and taxi and insurance companies as sellers), social media marketplaces, delivery service (e.g., connecting individuals to a myriad of businesses, like restaurants), digital investment (e.g., funds from different banks as products), and homestay platforms. It is noteworthy that many of these applications report data management challenges that we address in this work: event ordering and data invariants (Ferreira and Wible, 2017; Sarma, 2020; Yang et al., 2021; Singla and Wu, 2020; Chew, 2019; Silnitsky, 2022a, b) transactional consistency across microservices (Gluck, 2022; Khisti, 2018; Chornyi, 2018; Ferreira, 2019; Stenberg, 2019), consistent data querying  (Mayerhofer, 2018; Oliveira, 2017; Conger, 2021) data replication correctness (Viennot et al., 2015; Sarma, 2020; Andreakis et al., 2019).
@@ 3.3. Microservices
[Partial Data model] [Events]
@@
Traditionally, a microservice-based application is often made up of independent services, each deployed in a container on an OS-level virtualized platform such as Docker (Laigner et al., 2021). However, the emergence of programming frameworks for designing distributed applications, such as Orleans (Bykov et al., 2010) and Statefun (Statefun, 2023d), provides programming models that allow microservices to be implemented using higher-level abstractions, such as actors and functions, respectively.
To allow this benchmark to support the gamut of microservice implementations, we withdraw ourselves from defining an architectural blueprint, but rather, we focus on describing the expected independent components that must compose the benchmark application. In this sense, in the following paragraphs, we take advantage of Figure 1 to clarify the interactions, APIs, and the data model prescribed for each microservice. It is worth noting that we use the term ”events” as a communication abstraction for microservices, but these can also be framed as any message payload asynchronously delivered to a microservice. Besides, the event identifiers in Figure 1(b) do not imply a particular order in which the microservices exchange events. Furthermore, while we specify the microservices’ state and the queries using the relational model, they can also be specified through other data and query models as long as the same functionalities can be achieved.
Cart. The Cart microservice allows the customer to manage the products that ought to compose a checkout order. It does so through APIs for managing cart items and submitting cart checkouts.
To manage a customer’s cart, two relations are used: carts to track the status of a customer session, and $cart\_items$ to store the items that customers want to buy. Moreover, the $replica\_products$ relation represents a partial replica of the product table belonging to the Product microservice. The table is only updated through ProductUpdate events received from the Product microservice (represented by a dotted blue arrow in Figure 1(a)). The replication semantics are discussed in Section 4.3.
When a checkout request is received, Cart ensures the correctness of the customer’s cart by matching the cart items to the replicated product information and looking for products that had their price updated. In this case, the price is updated and the difference (if positive) between the new and the old price becomes a discount in the cart item. The customer’s cart items are then assembled into a ReserveInventory event and published for asynchronous processing. Once the event is published, the cart is sealed, and the customer is able to initiate a new cart session.
@@
Stock. The Stock microservice manages the inventory through a single relation: $stock\_item$. To update the inventory, the Stock microservice processes four events: ReserveInventory, PaymentConfirmed, PaymentFailed, and ProductDelete. Every inbound event leads to updates in one or more $stock\_item$’ tuples (depending on the number of items present in the checkout).
@@
Upon receiving a StockConfirmed event, the amount to charge the customer is calculated (including freight and discounts) and an invoice number is generated, resulting in a InvoiceIssued event. During this processing, all four relations have tuples created (in case it is the first customer order, otherwise $customer\_order$ tuple is updated). Afterwards, there are two moments where a customer’s order status is updated: after a payment and after shipment updates. Both trigger writes to order and $order\_history$ relations. In this case, the order of event processing here plays a role on maintaining the notion of time progress for individual orders.
Payment. The Payment microservice is responsible for managing payment data. Two relations are used to track an orders’ payment data: $payment$ and $card\_payment$. Payment processing starts through processing an InvoiceIssued event. Every credit applied to the order, namely, discounts and the payment option chosen by the customer at the time of the checkout (credit/debit card or a bank slip) are stored in $payment$ relations and in case of a card payment, a $card\_payment$ tuple is also inserted.
@@
The relation $order\_entry$ represents an order item but from the perspective of a seller. In this sense, amounts related to the item (i.e., discount, freight, total) are calculated. Every order (derived from the InvoiceIssued payload) is transformed into $N$ $order\_entry$ tuples, where N is the number of items corresponding to a seller. The relation $order\_entry$ is updated through processing the following events: PaymentConfirmed, PaymentFailed, ShipmentNotification, and DeliveryNotification.
@@ 3.4.1. Business Transactions
To realize the application scenario, we describe four business transactions that reflect different complexities in terms of the number of microservices involved and the number of events processed. In Section 4, we discuss how transactions possess different properties.
Customer Checkout. It starts in Cart microservice through processing a Checkout request. It involves Cart, Stock, Order, Payment, Shipment, Seller, and Customer microservices. A success path involves the following events: E2 - E3 - E5 - E6 - E8. There are two error cases: (i) when a payment fails, then from E5 we have E7 as the last; and when no single item from the cart is available in stock: E2 - E4. Although not shown in Figure 1 due to space constraints, a success path also involves E6 being processed by Order.
Price Update. To enable the partial replication of products in the Cart, upon processing a UpdatePrice request and updating its private state accordingly, the Product microservice generates E0 and sends it to the Cart microservice. By processing E0, the Cart microservice applies the new price to its corresponding replica.
Product Delete. To simulate making a product unavailable to customers, we pick a seller and a corresponding product (both from a distribution) and we set the product as disabled. To maintain the total number of products, thus avoiding anomalies in the distribution, we replace the deleted product with another one. This operation is done by processing a DeleteProduct request in Product microservice. Upon updating its private state accordingly, Product generates E1 and sends it to Stock.
Update Delivery. To simulate the delivery of goods, we pick the first 10 sellers with uncompleted orders (i.e., at least one package has not been delivered yet) in chronological order, and we set their respective oldest order’s packages as delivered. Thus, packages are progressively delivered as more update delivery transactions are submitted to the system. For maintenance of statistics about customers and orders, the following events are generated with the transaction: E9 is generated by Shipment for every package delivered and sent to the Seller and Customer microservices; and E8 is generated by Shipment when all packages of a shipment are delivered and sent to Order.